OpenStudy (anonymous):

Stan evaluated a function with a value and got 3 as the result. He wrote the entire process as a sentence. The function A.f(x)=x^2+2x B.f(x)=x^2-3x , When evaluated with A.3 B.-3 gives the value of 3 which function gices the value 3 when plugged in with 3 or -3 ?

OpenStudy (anonymous):

i do not understand what this question is asking let me try to read it again

OpenStudy (anonymous):

first off you have two functions

OpenStudy (anonymous):

yes function a and b

OpenStudy (anonymous):

not "a function" but if you get 3 for the first one, it means \[f(x)=x^2+2x=3\]

OpenStudy (anonymous):

oooh i see what it is asking

OpenStudy (anonymous):

plug in 3 in to each function and see which one gives 3 as an answer

OpenStudy (anonymous):

ive tried that but i keep getting 15 or -15

OpenStudy (anonymous):

for the first one?

OpenStudy (anonymous):

the question asks "which function gices the value 3 when plugged in with 3 or -3 ?"

OpenStudy (anonymous):

so evidently it is not A since \[3^2+2\times 3=15\]

OpenStudy (anonymous):

however if you try B you should get it

OpenStudy (anonymous):

for f(x)=x^2+2x i got -15 when plugged in with -3 and for f(x)=x^2-3x i got 0 when plugged in with -3

OpenStudy (anonymous):

yeah true but if you plug in 3 in to the second one, you should get 3 out

OpenStudy (anonymous):

btw the second one does not get 0 if you plug in 3

OpenStudy (anonymous):

hmmm hold on a second

OpenStudy (anonymous):

\[f(x)=x^2+2x\] \[f(-3)=(-3)^2+2\times (-3)\] try that

OpenStudy (anonymous):

you do not get \(-15\) because \((-3)^2=9\) not \(-9\) you should get \(9-6=3\) for A

OpenStudy (anonymous):

that was correct thank you so much can i ask you just one more question

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

A school is planning an event to raise $1,000 for charity. The cost of organizing the event is $200. The tickets are priced at $6 per child and $8 per adult. If x children and y adults attend the event, the equation that represents the total revenue that the organizers must collect from ticket sales to raise the targeted amount of $1,000 after paying for the $200 in expenses is . If the event were only for adults, adults would have to attend for the organizers to raise the targeted amount.

OpenStudy (anonymous):

how many adults would have to attend for the organizers to raise the targeted amount.? sorry

OpenStudy (anonymous):

& how would you set it up as a equation with the given information

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

can you please help me on this question

OpenStudy (anonymous):

@misty1212